OpenStudy (anonymous):

Prove by induction for n>= 1 1(2) + 2(3) + 3(4) +...+ n(n + 1) = (n(n + 1)(n + 2))/3

OpenStudy (turingtest):

step 1: prove it first for n=1

OpenStudy (turingtest):

have you done that?

OpenStudy (anonymous):

yes, I'm failing on the inductive step also proved it for n=2, n=3 for kicks

OpenStudy (turingtest):

ok, let me try it with you and we'll see how it goes...

OpenStudy (anonymous):

tried 1(2) + 2(3) + 3(4) + n(n + 1) + (n(n + 1))(n(n + 2)) =, does the left side seem right?

OpenStudy (turingtest):

so that is where your mistake probably lies...

OpenStudy (turingtest):

\[n=1:\]\[1(2)={1(1+1)(1+2)\over3}=2\]\[n:\]\[1(2) + 2(3) + 3(4) +...+ n(n + 1) ={ (n(n + 1)(n + 2))\over3}\]\[n+1:\]\[1(2)+2(3)+...+n(n+1)+(n+1)(n+2)\]\[={ (n(n + 1)(n + 2))\over3}+(n+1)(n+2)\]hope the algebra works out...

OpenStudy (anonymous):

got a math processing error, i'm guessing the error is in the +2 section

OpenStudy (turingtest):

you can't read above?

OpenStudy (anonymous):

came through error

OpenStudy (anonymous):

1(2) + 2(3) + 3(4) + n(n + 1) + (n+1)(n+2) = ?does the left side look right now?

OpenStudy (turingtest):

well I did 1(2)+2(3)+...+n(n+1)+(n+1)(n+2) yeah there we go

OpenStudy (anonymous):

cool, but the right side will end up cubic, and the left side quadratic

OpenStudy (turingtest):

you don't need to worry about the left, we just need to get ((n+1)(n + 2)(n + 3))/3 on the right and we will have proven it for n+1, and therefor all n

OpenStudy (experimentx):

summation of n is n(n+1)/2 summation of n^2 is n(n+1)(2n+1)/6 adding these we get, your required result

OpenStudy (turingtest):

@experimentX that is not the problem we are doing

OpenStudy (anonymous):

but don't we want the left side to equal the right side, and will the proof on the right side be (n(n + 1)(n + 2))/3 ?

OpenStudy (experimentx):

somewhat like this work?? http://www.wolframalpha.com/input/?i=sum+j%2Bj%5E2%2C+j%3D1+to+n

OpenStudy (turingtest):

no, we have to show that if we add one more term to the left, then the right side will give the same formula, but with n+1 in place of n

OpenStudy (anonymous):

so we don't even work with the right, just show that the right becomes one more than our hypothesis

OpenStudy (anonymous):

the algebra ends up being 2n^2 + 4n + 22 for n = 4, but now what?

OpenStudy (turingtest):

we are proving that the truth of the statement for n \(implies\) that it is true for n+1 sorry, but I don't think you understand the concept here jobu

OpenStudy (experimentx):

@TuringTest i mean it is same as the summation of j+j^2 for 1 to n, isn't it

OpenStudy (anonymous):

yes, turing, you understood it, but I'm stuck with what to do with the algebra after calculating the N+1

OpenStudy (turingtest):

we don't need to try various numbers for n @experimentX no it would be\[\sum_{i=1}^{n}i(i+1)\]

OpenStudy (turingtest):

oh yeah, sorry same thing

OpenStudy (turingtest):

I admit my error, but I still don't think that your solution is the intended method

OpenStudy (experimentx):

yeah thats what i mean, you can individually add up those terms, i and i^2

OpenStudy (turingtest):

gotchya, and I like your proof, but it is not by induction ;)

OpenStudy (anonymous):

so 1(2) + 2(3) + 3(4) + n(n + 1) + (n+1)(n+2) will somehow yield the desired result?

OpenStudy (anonymous):

meaning ((n+1)(n + 2)(n + 3))/3

OpenStudy (turingtest):

\[\sum_{i=1}^{n}i(i+1)={n(n+1)(n+2)\over3}\]we have to show that this implies that\[\sum_{i=1}^{n+1}i(i+1)={(n+1)((n+1)+1)((n+1)+2)\over3}\](i.e. we get the same formula as we do with n, but replaced with n+1)

OpenStudy (anonymous):

keep getting processing errors, i think I'm so close to understanding

OpenStudy (turingtest):

damn, forgot about the error thing, so look you are very close he have the sum up to n 1(2)+2(3)+...+n(n+1)=n(n+1)(n+2)/3 now the next term to both sides gives 1(2)+2(3)+...+n(n+1)+(n+1)(n+2)=n(n+1)(n+2)/3+(n+1)(n+2) the left side is clearly true because all we did was add the next term so we need to show through algebra that n(n+1)(n+2)/3+(n+1)(n+2)=(n+1)(n+2)(n+3)/3 and it will be proven

OpenStudy (turingtest):

we have*

OpenStudy (anonymous):

yessssss, i get it now, in essence, we don't need to worry about the LEFT side, but rather, add one to the right side

OpenStudy (turingtest):

exactly! glad you got the concept :)

OpenStudy (anonymous):

keep up the good work

OpenStudy (turingtest):

likewise!